### v1

@@ 4.4 Ablation studies
Figure 4: Visual comparison of ablation experiments. “CoSE” represents the result of the model where the decoder block is the CoSE module; “CoSE+CSI” represents the result of adding the CSI module to the previous model; “FIF-UNet” represents the result of Small FIF-UNet. The part of the white rectangular box is the place where there is an obvious segmentation error.
To confirm the effectiveness of the three proposed modules, the Synapse dataset is selected to conduct ablation experiments, in which we also consider the model performance by utilizing the pre-trained encoder. The experimental design for ablation studies concerns the employment of the proposed technical modules separately or their combinations with the baseline. The experimental results are listed in Table 3 in terms of concerned metrics. It can be found that all the proposed technical modules contribute to expected performance improvements. To be specific, by incorporating each module separately into the baseline model, the performance is improved by 0.9%, 0.89% and 0.88% in DICE score, respectively. In addition, the combinations of each two modules further enhance the segmentation performance with absolute improvements of 1.35%, 1.00% and 1.33%. Finally, the proposed FIF-UNet with all three modules harvests a 1.48% improvement over the baseline.
@@ 4.5 Generalization studies
In summary, the above experimental results show that all the proposed modules contribute to performance improvement even with different backbones, presenting expected generalization ability with consistent performance. Most importantly, we can also observe that the proposed modules have the ability to obtain higher performance improvements for simpler model architecture (i.e., original UNet).
4.6 Visualization results
@@
To better understand the proposed model, visualization results of ablation studies are provided in Figure 4. According to the results in Table 3, for the single module and the combination of two modules, the models with the highest DICE are selected for visualization analysis. Compared with the baseline model, the CoSE in the decoder considerably enhances the model to recognize foreground and background features, resulting in significantly reduced error regions. In particular, for the third row, the model successfully segmented the gallbladder objects, while the baseline model fails to identify them. With the combination of the CSI module, the approximate shape of each organ can be localized, suggesting that the model can accurately understand the structural characteristics of organs by obtaining informative features based on the different levels of semantic information between the encoder and the decoder. Compared with the previous model, the FIF-UNet has the ability to accurately capture the spatial distribution and morphological details of each organ by integrating semantic information of different scales, benefiting to reduce the misidentified scattering areas in the segmentation results.
@@
In addition, the classical SwinUNet and three recent models (PVT-CASCADE, TransCASCADE, and Cascaded MERIT) are also selected to compare the visualization results with the proposed FIF-UNet. As shown in Figure 5, compared with the selected models, the proposed FIF-UNet is capable of accurately locating all organs with precise details. For the segmentation of the stomach, the selected baselines suffer from task challenges due to its high similarity with the background, i.e., can only partially identify the organ, or even fail to detect them. Fortunately, the proposed FIF-UNet model demonstrates significant advantages in accurately segmenting the entire stomach and greatly reducing the risk of misidentifying other regions as the stomach. Furthermore, in the case of the second row, it can also be observed that compared to recent models, only the FIF-UNet can identify the hollow region in the middle of the left kidney, illustrating its advantages in capturing complex structural details. From the above analysis, except for the quantitative metrics, more qualitative results also support the performance improvements over baselines.
@@ 5 Conclusion
In this work, a novel model called FIF-UNet is proposed to achieve accurate medical image segmentation, which effectively leverages the semantic information of both the encoder and decoder stages by feature interaction and fusion. Three modules, CoSE, CSI and MLF, are designed for the FIF-UNet to enhance the skip connections and decoder. Experimental results on two public datasets demonstrate that the proposed model outperforms SOTA methods in terms of certain metrics. In addition, the proposed three modules contribute to desired performance improvements. Most importantly, they can also be adapted to other similar architectures to improve the overall performance, indicating the desired generalization ability in the segmentation tasks.

### v2

@@ 4.4 Ablation studies
To comprehensively evaluate the effectiveness of each proposed component, we conduct a series of ablation studies on two benchmark datasets. Specifically, we first analyze the contribution of each module (CSI, CoSE, and MLF) on the Synapse dataset, followed by a similar investigation on the ACDC dataset to verify the generalization ability across different tasks. In addition, we compare different loss functions to assess their impact on segmentation performance. Finally, we evaluate the superiority of our CSI module by comparing it with alternative skip connection enhancement strategies.
4.4.1 Ablation experiment on Synapse dataset
@@
To validate the effectiveness and complementarity of each proposed module, we conducted ablation studies on the Synapse dataset, as summarized in Table 4. The experimental design for ablation studies concerns the employment of the proposed technical modules separately or their combinations with the baseline.
It can be seen that the CoSE module enhances the model’s ability to capture discriminative features, which is particularly beneficial for segmenting organs with low contrast to the background or ambiguous anatomical boundaries. Specifically, the CoSE improves the Dice scores for the SM and PC by 1.30% and 2.44%, respectively. The CSI module facilitates semantic fusion between encoder and decoder features to effectively preserve more contextual and detailed information, achieving improvements for both small and ambiguous organs, with gains of 3.44% and 2.62% on KL and KR, and 1.74% on the SM. The MLF module strengthens fine-grained feature representations through multi-level fusion to support the segmentation of small-scale structures, harvesting the most significant performance improvements for KL and KR by 3.93% and 2.99%, respectively.
In addition, by combining two of them into the proposed model, distinct advantages are observed depending on organ characteristics. To be specific, the combination of CSI and CoSE yields notable improvements for challenging organs such as the GB and PC, with Dice score increases of 0.96% and 2.42%, respectively. These results suggest that CSI’s semantic enhancement and CoSE’s discriminative focus are complementary. On the other hand, CSI+MLF and CoSE+MLF combinations show greater improvements for small organs, as they can preserve and refine detailed information at different stages of the network.
Finally, the full FIF-UNet model, incorporating all three modules, achieves the best overall performance with an average Dice score of 86.05%, representing a 1.48% improvement over the baseline. It is believed that the comprehensive enhancement across small, ambiguous, and difficult organs can validate the proposed modules and the overall effectiveness of the proposed architecture.
4.4.2 Ablation experiment on ACDC dataset
Table 5: Ablation studies based on the ACDC dataset
CoSE CSI MLF Avg DICE↑ RV Myo LV
$\times$ $\times$ $\times$ 91.94 89.96 90.01 95.84
$\times$ $\times$ 92.27 90.83 90.02 95.95
$\times$ $\times$ 92.30 90.79 90.25 95.86
$\times$ $\times$ 92.26 90.70 89.93 96.15
$\times$ 92.47 90.54 90.47 96.39
$\times$ 92.46 91.13 90.22 96.02
$\times$ 92.51 91.07 90.32 96.13
92.58 91.30 90.24 96.19
Table 5 presents ablation results on the ACDC dataset, evaluating the effectiveness and complementarity of the proposed CoSE, CSI, and MLF modules in cardiac structure segmentation. As can be seen from the results, each proposed module, whether independent employment or in combination, can also contribute desired performance improvements in this work.
For the individual modules, the CoSE module results in the most notable improvement in the segmentation of the RV (the Dice score increases by 0.87%), which confirms that CoSE enhances the model’s ability to capture discriminative features, particularly in regions with low contrast and ambiguous boundaries. The CSI module, designed to facilitate semantic interactions between encoder and decoder features, consistently improves performance across all three anatomical structures, increasing the average Dice score by 0.36%, which highlights CSI’s effectiveness in preserving and transmitting semantic information. The MLF module, which performs multi-scale feature fusion at the output stage, significantly boosts segmentation performance for the LV with a relatively regular shape and well-defined boundaries, which validates MLF’s strength in enhancing structural detail.
Similarly, further performance gains are observed for evaluating pairwise combinations of the modules. The integration of CSI and CoSE produces significant improvements for challenging structures such as the RV and Myo, with Dice scores reaching 90.54% and 90.47%, respectively, and an average Dice gain of 0.53%. The results suggest that combining discriminative feature enhancement with multi-level semantic interactions is particularly effective for segmenting complex or ambiguous anatomical regions. Similarly, the combination of CoSE and MLF improves RV and Myo segmentation with an average Dice gain of 0.52%, indicating that enhancing boundary details further strengthens discriminative feature representations. Although the CSI and MLF combination lacks the explicit discriminative modeling provided by CoSE, it still achieves the highest average Dice improvement (0.57%) among the pairwise configurations, due to effective integration of semantic context and structural detail at different stages of the network.
The full model FIF-UNet with all three modules achieves the best performance (average Dice: 92.58%, ↑ 0.64%), confirming their complementary roles and strong synergy in improving segmentation accuracy and robustness across all cardiac structures.
4.4.3 Comparison of different loss functions
Table 6: Comparative analysis of model performance based on different loss functions on Synapse dataset for organ segmentation. Organ abbreviations: GB (gallbladder), KL (left kidney), KR (right kidney), PC (pancreas), SP (spleen), SM (stomach). Only Dice scores are reported for individual organs. High Dice scores and low HD95 scores mean better performance. The best result is highlighted in bold, and the second-best is highlighted with an underline.
loss1 loss2 Average Aorta GB KL KR Liver PC SP SM
DICE↑ HD95↓
Generalized DICE CE Loss 84.36 15.24 89.76 69.21 88.77 84.69 95.37 72.26 90.31 84.54
Tversky Loss CE Loss 85.35 24.88 88.56 76.78 86.48 84.09 95.50 74.82 90.83 85.77
Dice Loss Focal Loss 85.42 17.28 88.60 76.36 87.62 84.86 95.29 74.11 91.08 85.42
Dice Loss CE Loss 86.05 15.82 89.49 76.15 88.23 86.26 95.87 74.14 91.31 86.97
Figure 5: Comparison of training loss and Dice score curves under different loss function combinations. Dice + CE Loss achieves the most stable convergence and the highest segmentation performance.
To evaluate the impact of different loss functions on segmentation performance, we perform a comparative study on the Synapse dataset, as shown in Table 6. Among all configurations, the combination of Dice Loss and CE Loss achieves the best overall performance, yielding the highest average Dice score (86.05%) and the second-lowest HD95 distance (15.82). In addition, this combination produces either the best or second-best Dice scores for most individual organs, indicating both high segmentation accuracy and strong anatomical consistency, which highlights the effectiveness and robustness of this loss formulation for medical image segmentation tasks.
The superior performance of the Dice + CE combination benefits from the complementary properties of the two loss functions. Dice Loss effectively addresses class imbalance by directly optimizing region overlap, which is especially beneficial for small or underrepresented organs. However, it often produces unstable gradients during early training. In contrast, Cross-Entropy Loss provides stable pixel-wise supervision, enables faster convergence, and improves boundary delineation in the initial stages. By combining them to perform the model training, CE Loss guides the model in capturing coarse structural features, while Dice Loss enhances global region alignment, resulting in improved segmentation accuracy and anatomical consistency.
Other loss function combinations offer certain advantages in specific scenarios but fail to achieve comparable overall performance. Generalized Dice Loss, although more suitable for extreme class imbalance, introduces additional complexity and typically yields lower Dice scores across most organs. Tversky Loss, designed to balance false positives and false negatives, depends heavily on task-specific tuning and fails to generalize across diverse classes. Focal Loss focuses on hard-to-classify pixels but lacks the region-level structural sensitivity provided by CE Loss and further leads to reduced robustness. These comparisons further confirm the effectiveness and generalization capability of the Dice + CE combination in multi-organ segmentation tasks.
To further investigate the impact of different loss function combinations on training loss and segmentation performance, the training convergence is shown in Figure 5. Among all configurations, the combination of Dice Loss and CE Loss exhibits the most stable and efficient training. The average loss decreases rapidly with minimal fluctuations and stabilizes early, while the Dice score reaches 0.8 within the first 50 epochs and remains consistently high results throughout the training process. In contrast, the Dice + Focal Loss demonstrates a similar initial convergence pattern but suffers from greater fluctuations in both the loss and Dice score, i.e., unstable performance.
The Tversky Loss + CE Loss combination results in poor convergence efficiency with unstable loss values. Although the Dice score increases relatively fast in the early stages, it can only obtain inferior final results with limited overall learning capacity. The Generalized Dice + CE Loss shows the fastest decline in loss, prematurely approaching zero and further to overfitting. Meanwhile, the resulting Dice score improves slowly and remains the lowest among all configurations. In summary, these observations confirm that the Dice + CE Loss combination achieves the best balance between fast convergence, stable optimization, and final segmentation accuracy.
4.4.4 Comparison of different skip connection enhancement methods
Table 7: Comparative analysis of model performance based on different skip connection enhancement methods on Synapse dataset for organ segmentation. Organ abbreviations: GB (gallbladder), KL (left kidney), KR (right kidney), PC (pancreas), SP (spleen), SM (stomach). Only Dice scores are reported for individual organs. High Dice scores and low HD95 scores mean better performance. The best result is highlighted in bold, and the second-best is highlighted with an underline.
Methods Average Aorta GB KL KR Liver PC SP SM
DICE↑ HD95↓
AG 84.06 17.78 89.05 71.21 86.56 83.29 94.96 73.92 90.74 82.77
FFM 85.15 18.77 88.88 73.74 85.11 84.16 95.40 73.71 92.96 87.24
GAB 85.16 15.83 88.42 74.62 89.56 84.91 94.99 72.02 91.71 85.06
concatenation 85.39 11.98 89.44 75.90 91.25 86.55 95.67 74.07 87.41 82.84
SCCSA 85.57 15.65 89.33 79.36 88.66 85.65 95.13 73.55 90.97 81.94
Addition 85.57 14.08 88.87 74.47 88.71 85.36 95.40 73.57 93.02 85.13
CSI(ours) 86.05 15.82 89.49 76.15 88.23 86.26 95.87 74.14 91.31 86.97
To evaluate the superiority of the proposed CSI module in enhancing skip connections, we conduct a comparative study over several representative fusion strategies, including AG (Attention Gate in Attention U-Net (2018)), FFM (Feature Fusion Module in FusionU-Net (2024)), GAB (Group Aggregation Bridge in EGE-UNet (2023)), and SCCSA (Skip Connection Channel-Spatial Attention in BRAU-Net++ (2024)), as well as element-wise addition and channel concatenation. In each case, only the CSI module in FIF-UNet is replaced with the alternative module to conduct the performance comparison.
As shown in Table 7, the proposed CSI module achieves the highest average Dice score of 86.05%, outperforming all baselines with selected fusion modules. In terms of organ-wise performance, CSI achieves the best Dice scores on the aorta (89.49%), liver (95.87%), and pancreas (74.14%), and ranks second-best on the gallbladder, right kidney, and stomach, showing strong generalization across organs with varying sizes and complexities. Compared to the best baseline (SCCSA and Addition, both 85.57%), the CSI enables the model to obtain a 0.48% improvement, demonstrating the superiority of the fusion design. In terms of HD95, Concatenation achieves the lowest value (11.98) due to its simple channel-stacking operation with fine-grained details, especially for small organs like kidneys, while it only obtains undesired overall performance due to limited semantic filtering.
The AG module applies decoder features to reweight encoder features via gating mechanism, but the one-way gating strategy may discard valuable decoder information with reduced representation ability. The FFM fuses multi-scale encoder features without decoder features entirely, ignoring crucial context for semantic alignment. The GAB leverages decoder-generated masks to guide fusion of encoder features, in which, as decoding progresses, the resolution of these masks becomes coarser and noise-sensitive, to further diminish the model’s effectiveness. The SCCSA has a similar structure to the proposed model, i.e., cascading channel and spatial attentions. However, the SCCSA first performs the addition operation for encoder and decoder features with a follow-up attention, which may limit its ability to model cross-level dependencies effectively. In contrast, the proposed CSI explicitly learns channel and spatial correlations between encoder and decoder features before performing element-wise addition, enabling more informative and meaningful semantic fusion.
In summary, the results confirm that the CSI effectively facilitates multi-level semantic interaction between encoder and decoder, enhancing the representation ability of skip connections and benefiting to superior multi-organ segmentation performance over existing fusion strategies.
@@ 4.5 Generalization studies
In summary, the above experimental results show that all the proposed modules contribute to performance improvement, even with different backbones, presenting expected generalization ability with consistent performance. Most importantly, we can also observe that the proposed modules have the ability to obtain higher performance improvements for simpler model architectures (i.e., original UNet).
4.6 Qualitative results
To further validate the quantitative findings and demonstrate the effectiveness of both the proposed modules and the overall FIF-UNet architecture, we provide comprehensive visualizations on the Synapse and ACDC. For each dataset, the ablation visualizations are based on the models that achieved the highest Dice scores among the single-module and two-module combinations,
4.6.1 Experimental results on Synapse dataset
Figure 6: Visual comparison of ablation experiments on the Synapse dataset. “CoSE” represents the result of the model where the decoder block is the CoSE module; “CoSE+CSI” represents the result of adding the CSI module to the previous model; “FIF-UNet” represents the result of Small FIF-UNet. The part of the white rectangular box is the place where there is an obvious segmentation error.
@@ 4.6.1 Experimental results on Synapse dataset
Visualization results of ablation studies on the Synapse dataset are provided in Figure 6. Compared with the baseline model, the CoSE in the decoder considerably enhances the model to recognize foreground and background features, resulting in significantly reduced error regions. In particular, for the third row, the model successfully segments the gallbladder objects, while the baseline model fails to identify them. With the combination of the CSI module, the approximate shape of each organ can be localized, suggesting that the model can accurately understand the structural characteristics of organs by obtaining informative features based on the different levels of semantic information between the encoder and the decoder. Compared with the previous model, the FIF-UNet has the ability to accurately capture the spatial distribution and morphological details of each organ by integrating semantic information of different scales, which helps reduce the misidentified scattering areas in the segmentation results.
@@
In addition, the classical SwinUNet and three recent models (PVT-CASCADE, TransCASCADE, and Cascaded MERIT) are also selected to compare the visualization results with the proposed FIF-UNet. As shown in Figure 7, compared with the selected models, the proposed FIF-UNet is capable of accurately locating all organs with precise details. For the segmentation of the stomach, the selected baselines suffer from task challenges due to its high similarity with the background, i.e., can only partially identify the organ, or even fail to detect it. Fortunately, the proposed FIF-UNet model demonstrates significant advantages in accurately segmenting the entire stomach and greatly reducing the risk of misidentifying other regions as the stomach. Furthermore, in the case of the second row, it can also be observed that compared to recent models, only the FIF-UNet can identify the hollow region in the middle of the left kidney, illustrating its advantages in capturing complex structural details. From the above analysis, except for the quantitative metrics, more qualitative results also support the performance improvements over baselines.
4.6.2 Experimental results on ACDC dataset
Figure 8: Visual comparison of ablation experiments on the ACDC dataset. “CSI” represents the result of the model where the skip connection is the CSI module; “CSI+MLF” represents the result of adding the MLF module to the previous model; “FIF-UNet” represents the result of Small FIF-UNet. The part of the blue rectangular box is the place where there is an obvious segmentation error.
Figure 9: Visual comparison of different models on the ACDC dataset. “FIF-UNet” denotes Small FIF-UNet. The part of the blue rectangular box is the place where there is an obvious segmentation error.
The visualization results of the ablation experiments on the ACDC dataset are illustrated in Figure 8. In general, the CSI module enhances the model’s ability to capture global semantic information to further improve background suppression and structure delineation. As seen in the first row, the CSI can correct background misclassification from the baseline model to highlight its semantic enrichment. However, in ambiguous cases with high organ-background similarity, such as the RV shown in the third row, only the CSI fails to accurately localize the RV. As in the second row, by adding the MLF module, the proposed model improves fine-grained detail through multi-scale feature fusion, i.e., the RV boundary is better recovered to indicate enhanced detail sensitivity. The full FIF-UNet model, incorporating CoSE, CSI and MLF, achieves the most robust segmentation results. The CoSE strengthens the model’s discriminative ability via channel-wise attention, enabling accurate identification of visually similar regions. In the third row, only FIF-UNet can successfully segment the RV, demonstrating the critical role of CoSE in refining feature representation and improving segmentation in challenging scenarios.
Figure 9 presents qualitative comparisons between FIF-UNet and four methods: SwinUNet, TransCASCADE, Cascaded MERIT, and Parallel MERIT. Similar to the results on the Synapse dataset, FIF-UNet consistently shows superior segmentation accuracy, particularly for organs with low contrast or indistinct boundaries. In the first row, SwinUNet and TransCASCADE fail to segment the RV, while Cascaded MERIT and Parallel MERIT only partially recover the structure. In contrast, FIF-UNet successfully identifies the entire RV, indicating its enhanced robustness in segmenting low-contrast, ambiguous structures. The second row features a highly blurred RV boundary, which provides considerable challenges to all comparative models. Even under these conditions, FIF-UNet accurately recovers more of the RV region than its counterparts, further demonstrating its strength in handling edge ambiguity. Finally, the third row shows that FIF-UNet produces the fewest false positives among all models, showing its superior performance and reduced susceptibility to misclassification.
@@ 5 Conclusion
In this work, we propose FIF-UNet, a novel U-shaped architecture designed to improve medical image segmentation by fully leveraging multi-level semantic features through effective interaction and fusion. The model integrates three plug-and-play modules: the Channel-Spatial Interaction (CSI) module, which enhances skip connections by modeling cross-stage dependencies from both channel and spatial perspectives; the Cascaded Convolution and Squeeze-and-Excitation (CoSE) module, which enhances the decoder’s ability to focus on critical regions by incorporating channel attention; and the Multi-Level Fusion (MLF) module, which aggregates decoder features across scales to preserve structural details. Extensive experiments on the Synapse and ACDC datasets demonstrate that FIF-UNet consistently outperforms state-of-the-art methods. Furthermore, ablation studies confirm the individual contributions of each module. Notably, all proposed modules are generalizable, making them applicable to other encoder-decoder segmentation frameworks.